# Universal Adaptive Constraint Propagation: Scaling Structured Inference for Large Language Models via Meta-Reinforcement Learning

### v1

@@ Abstract
Universal Adaptive Constraint Propagation: Scaling Structured Inference for Large Language Models via Meta-Reinforcement Learning
@@ 1 Introduction
Large language models increasingly require structured inference where outputs must satisfy complex constraints: JSON schema enforcement for API calls, formal logic generation for reasoning tasks, and multi-lingual parsing for cross-lingual applications. Consider a modern LLM generating a structured response that must conform to a JSON schema with nested objects, required fields, and type constraints. The order in which these constraints are checked and propagated determines both generation speed and correctness. Current LLM inference pipelines rely on post-hoc validation or static constraint checking, leading to wasted computation when invalid continuations are generated early. Similarly, in multi-lingual parsing, the same propagation principles should apply across languages, yet existing systems require language-specific tuning. The challenge is to learn universal constraint propagation policies that generalize across languages, tasks, and constraint types without expensive per-task retraining.
In many of these settings, inference can be naturally cast as constraint propagation over an internal state that encodes partial hypotheses, variable domains, and active constraints. The order in which these constraints are processed, the propagation schedule, has a strong effect on both runtime and solution quality. Static scheduling policies such as First-In-First-Out queues or simple degree-based heuristics are widely used, but they are myopic: they do not account for how a local propagation step will affect the future evolution of the inference state, and they cannot adapt to the structure of a particular instance or domain.
Reinforcement learning offers a way to turn propagation scheduling into a sequential decision problem. Recent advances in deep RL have demonstrated that learned policies can outperform handcrafted heuristics in complex domains such as games and combinatorial optimization (silver2016alphago; silver2017alphago; mnih2015human; lillicrap2015continuous; watkins1992q; williams1992simple). By treating each propagation decision as an action taken in response to the current inference state, we can train an agent to select constraints that maximize long-term gains, such as overall pruning or final task accuracy, rather than relying on static local rules (lecun1998gradient; hochreiter1997long; rumelhart1986learning).
@@
In this paper, we introduce MetaJuLS, a meta-reinforcement learning approach for universal adaptive constraint propagation that scales to modern LLM inference. We formulate structured inference as constraint propagation where variables represent linguistic or structural units, domains encode candidate outputs, and constraints enforce grammar rules, schemas, or logical requirements. The key innovation is a meta-learned policy parameterized by a Graph Attention Network (GAT) (velickovic2018graph) that adapts to new grammars, languages, or constraint types in seconds via few-shot fine-tuning, rather than requiring hours of task-specific training. We train the policy using Model-Agnostic Meta-Learning (MAML) (finn2017model) over diverse constraint propagation tasks, enabling rapid adaptation across languages and domains. We reserve “zero-shot” for transfer requiring no gradient steps; our setting is few-step adaptation (5–10 gradient steps, 5–15 seconds).
Although constraint programming is traditionally studied outside NLP, many language inference problems reduce to constraint-based reasoning, making CP benchmarks a complementary validation rather than the focus. We therefore also evaluate MetaJuLS on classical constraint programming benchmarks drawn from MiniZinc and XCSP, not to claim CP state-of-the-art, but to test whether the scheduling principles learned for language inference extend to other constraint-based reasoning tasks.
Our contributions are as follows:
@@
We formulate a unified constraint-propagation scheduling framework spanning constituency parsing, dependency parsing, and LLM constrained decoding, demonstrating that learned scheduling policies achieve 1.5–2.0 speedups over GPU-optimized baselines while maintaining within 0.2% accuracy (Section 4).
We introduce meta-learning for rapid adaptation across languages and tasks, enabling few-step adaptation (5–10 gradient steps, 5–15 seconds) rather than hours of task-specific training (Section 4.5).
@@
We demonstrate that MetaJuLS scales to modern LLM inference, accelerating JSON schema enforcement and formal logic generation (LogicBench, GSM8K-Constrained) by 1.8 compared to speculative decoding, with factorial experiments showing orthogonality (Section 4.3).
@@
We provide ablation-backed evidence that RL is necessary: MetaJuLS outperforms strong heuristic baselines (VSIDS-style, cost-normalized greedy, supervised ranking) by 1.2–1.5 in speed while matching or exceeding accuracy (Section 4).
@@
We introduce a safety-aware fallback mechanism with principled entropy-based guarantees, reducing accuracy gaps from 2.1% to 0.15% while maintaining 1.4 average speedups (Section 4.4).
@@ 2.1 Constituency Parsing as Constraint Propagation
Constraint propagation then proceeds by repeatedly selecting a constraint whose scope intersects with recently modified variables, applying a propagator that tightens the domains of those variables, and marking downstream constraints as potentially affected. This formulation inverts the typical constraint propagation framing: rather than pruning invalid values, parsing constraints derive valid constituents. We retain the term “propagation” because the computational structure is identical: constraints are activated by domain changes and processed according to a schedule, and scheduling remains the key efficiency lever. Under beam search or pruned inference, propagation order determines which constituents are derived before the beam fills, directly impacting both coverage and speed. When domains are interpreted as sets of possible labels or structures for linguistic units, this propagation view aligns closely with existing dynamic programming and message-passing algorithms for NLP, but makes explicit that the order in which constraints are processed is a policy choice rather than a fixed artifact of a particular algorithm.
@@ 2.3 Related Work
Activity-based heuristics, inspired by the VSIDS (Variable State Independent Decaying Sum) strategy from SAT solving (moskewicz2001chaff; marques2012activity), represent a significant advance over static scheduling (regin1994filtering; lecoutre2009str). These methods maintain activity scores for constraints based on their historical contribution to search progress, dynamically prioritizing constraints that have recently caused domain reductions or failures (audemard2013glucose; marques2009glucose; debruyne1997optimal). However, activity-based heuristics remain reactive rather than predictive, adjusting priorities only after observing constraint behavior rather than anticipating which constraints will be most effective in the current context.
Our work builds on these foundations by combining graph neural networks with reinforcement learning to learn adaptive propagation policies (velickovic2018graph; schulman2017proximal). Unlike previous approaches that rely on handcrafted features or reactive heuristics (wallace1996practical; marques2012activity), MetaJuLS learns to extract relevant information from the constraint-variable graph structure itself (kipf2017semi; hamilton2017inductive), enabling the discovery of novel scheduling strategies that adapt to both problem structure and solver state (bessiere2006constraint; van2006propagation).
@@ 3.1 Problem Formulation
The action space consists of all constraint propagators that are currently “dirty,” that is, constraints whose associated variables have experienced domain changes since their last propagation (bessiere2006constraint; van2006propagation). At each time step , the agent selects an action , where is the subset of dirty constraints at time . The selected propagator is then executed, potentially reducing variable domains and marking additional constraints as dirty for future propagation (regin1994filtering; lecoutre2009str).
@@
where is the set of variables involved in constraint , is the current domain size of variable , and is a constraint-type-specific constant reflecting the complexity of the filtering algorithm. The constants reflect empirical operation counts: grammar rule constraints require checking two child domains (binary lookup), while lexical constraints require a single vocabulary check. These values are robust to perturbation; Appendix A.1 reports 2% performance variation for . For grammar rule constraints in parsing, we set (binary checks), while for lexical constraints (simpler lookups). For CP constraints, for binary constraints and for alldiff where is arity. This deterministic proxy avoids hardware-dependent wall-clock time measurements while accurately reflecting computational cost (schulman2017proximal). To validate the cost proxy, we measure correlation between proxy cost and wall-clock time across constraint types and sentence lengths. On PTB parsing, we find Pearson correlation (, ) between proxy cost and measured microseconds per propagator, confirming that optimizing the proxy improves real wall-clock time. The reward structure encourages the agent to discover policies that maximize search space reduction while minimizing solver overhead (debruyne1997optimal; gent2006watched).
@@ 4.1 Constituency Parsing: Penn Treebank
We compare against Berkeley Neural Parser (kitaev2018constituency), GPU-parallelized CKY (our reimplementation of klein2003accurate using batched CUDA kernels for span-parallel chart filling), A* search with neural heuristics (stern2017minimum), activity-based scheduling, VSIDS-style conflict-driven scheduling (prioritizing constraints that recently caused pruning), cost-normalized greedy (maximizing using our deterministic cost proxy), and supervised ranking (learning-to-rank constraints from state features, trained to imitate oracle best-first on development set). The meta-learned MetaJuLS policy is trained on PTB sections 2–21. Table 1 reports F1 scores and inference time. MetaJuLS achieves 1.6–1.9 speedups over GPU-parallelized CKY while maintaining within 0.2% F1 of Berkeley Neural Parser.
@@ 4.2 Dependency Parsing: Universal Dependencies
We evaluate MetaJuLS on dependency parsing using Universal Dependencies (UD) v2.11 (nivre2016universal) across 10 diverse languages: English, Spanish, French, German, Chinese, Arabic, Finnish, Japanese, Russian, and Hindi. We implement a transition-based dependency parser where variables correspond to parser states, domains encode possible transitions, and constraints enforce arc consistency and well-formedness. All UD comparisons use the same dependency parser architecture, beam size, and batching; only the constraint scheduling policy changes, ensuring apples-to-apples comparison.
We compare against state-of-the-art dependency parsers (UDPipe (straka2016udpipe), Stanza (qi2020stanza)), activity-based scheduling, VSIDS-style conflict-driven scheduling, cost-normalized greedy, and MetaJuLS. The meta-learned MetaJuLS policy is trained on English UD data, then rapidly adapted to other languages with 5–10 gradient steps (requiring 5–15 seconds). Table 2 reports LAS (Labeled Attachment Score) and inference time. MetaJuLS achieves 1.5–1.8 speedups over activity-based scheduling while maintaining competitive LAS scores.
@@ 4.3 LLM Constrained Decoding
We evaluate on LogicBench (chen2024logicbench), which requires generating formal logical expressions that satisfy given premises. Variables correspond to token positions, domains contain logical operators and predicates, and constraints enforce well-formedness and logical consistency. To test whether MetaJuLS and speculative decoding are complementary, we perform a factorial experiment comparing: (1) baseline (neither method), (2) speculative decoding only, (3) MetaJuLS scheduling only, and (4) both methods combined. We compare against NeuroLogic Decoding (lu2021neurologic), Outlines, and LMQL. MetaJuLS achieves 1.8 speedup over speculative decoding alone while maintaining 98.2% constraint satisfaction rate (vs. 94.1% for speculative decoding). When combined, speculative decoding + MetaJuLS achieves an additional 15% speedup over MetaJuLS alone, demonstrating orthogonality.
@@ 4.5 Rapid Cross-Lingual and Cross-Task Adaptation
4.6 Mechanistic Interpretability: Probing Studies

### v2

# Adaptive Constraint Propagation: Scaling Structured Inference for Large Language Models via Meta-Reinforcement Learning

@@ Abstract
Adaptive Constraint Propagation: Scaling Structured Inference for Large Language Models via Meta-Reinforcement Learning
@@ 1 Introduction
Large language models increasingly operate under hard output constraints: JSON schemas for tool/API calls, well-formed logical forms for reasoning, and syntactic constraints for multilingual parsing. In these settings, inference is not just scoring tokens—it is maintaining satisfiable partial structures. When constraint checks are applied late (post-hoc validation) or in a fixed order, models waste computation exploring invalid continuations and often require task- or language-specific tuning to remain robust.
A unifying view is to cast structured inference as constraint propagation over an evolving state of partial hypotheses, variable domains, and active constraints. The key efficiency lever is the propagation schedule: which constraint to process next among those made relevant by recent state changes. Standard schedulers—FIFO queues, degree-based ordering, or reactive activity heuristics—are inexpensive but myopic, because they optimize local effects rather than anticipating downstream pruning and search dynamics.
This naturally suggests a sequential decision problem: choose the next propagation to maximize long-horizon gains in pruning and solution quality under compute budgets. Reinforcement learning has repeatedly shown that learned policies can outperform handcrafted heuristics in domains where actions have delayed, global consequences (mnih2015human; silver2016alphago; silver2017alphago; lillicrap2015continuous; watkins1992q; williams1992simple). Here, each propagation step becomes an action conditioned on the current solver state, enabling the policy to trade off immediate cost against future reductions in search and error (rumelhart1986learning; lecun1998gradient; hochreiter1997long).
@@
We introduce MetaJuLS, a meta-reinforcement learning approach for transferable adaptive constraint propagation that scales to modern LLM inference. We represent the inference state as a constraint–variable graph and parameterize the scheduler with a Graph Attention Network (velickovic2018graph). To generalize across languages and constraint types, we meta-train the policy with Model-Agnostic Meta-Learning (finn2017model), enabling few-step adaptation (5–10 gradient steps; 5–15 seconds) rather than hours of task-specific retraining. We use “zero-shot” strictly for transfer with no gradient steps; our setting is rapid few-shot adaptation.
While constraint programming (CP) is not the paper’s focus, many NLP and LLM constraint systems share the same propagation mechanics. We therefore use MiniZinc and XCSP benchmarks as complementary stress tests to evaluate whether the learned scheduling principles transfer beyond linguistic constraints.
Our contributions are:
@@
A unified constraint-propagation scheduling formulation spanning constituency parsing, dependency parsing, and LLM constrained decoding, yielding 1.5–2.0 speedups over GPU-optimized baselines while staying within 0.2% of state-of-the-art accuracy (Section 4).
A meta-learning framework for rapid cross-language and cross-task adaptation, requiring only 5–10 gradient steps (5–15 seconds) instead of hours of retraining (Section 4.5).
@@
An LLM deployment result: faster JSON schema enforcement and formal logic generation (LogicBench, GSM8K-Constrained), including factorial experiments showing orthogonality with speculative decoding (Section 4.3).
@@
Ablation-backed evidence that RL is necessary: MetaJuLS outperforms strong heuristic and learned baselines (VSIDS-style, cost-normalized greedy, supervised ranking) by 1.2–1.5 in speed while matching or exceeding accuracy (Section 4).
@@
A safety-aware fallback mechanism using policy entropy to preserve accuracy, reducing gaps from 2.1% to 0.15% while maintaining 1.4 average speedups (Section 4.4).
@@ 2.1 Constituency Parsing as Constraint Propagation
Constraint propagation then proceeds by repeatedly selecting a constraint whose scope intersects with recently modified variables, applying a propagator that tightens the domains of those variables, and marking downstream constraints as potentially affected. This formulation inverts the typical constraint propagation framing: rather than pruning invalid values, parsing constraints derive valid constituents. We retain the term “propagation” because the computational structure is identical: constraints are activated by domain changes and processed according to a schedule, and scheduling remains the key efficiency lever. Under beam search or pruned inference, propagation order determines which constituents are derived before the beam fills, directly impacting both coverage and speed. When domains are interpreted as sets of possible labels or structures for linguistic units, this propagation view aligns closely with existing dynamic programming and message-passing algorithms for NLP, but makes explicit that the order in which constraints are processed is a policy choice rather than a fixed artifact of a particular algorithm (see Figure 1).
@@ 2.2 Why RL Is Necessary for Language Inference
In language, constraints often interact over long ranges and across multiple levels of representation. Subject-verb agreement couples distant tokens via syntactic structure; coreference constraints link mentions far apart in the surface string; semantic parsing constraints couple local predicate arguments with global type information. Static scheduling strategies cannot adapt to these context-dependent trade-offs, treating all propagation opportunities as roughly equivalent. Activity-based heuristics provide some adaptivity by reacting to past conflicts, but they are fundamentally reactive rather than predictive. Modern neural parsers operate under computational constraints such as beam limits, early stopping, and timeout budgets, where propagation order determines which hypotheses are explored before resources are exhausted, making scheduling a first-order concern for practical deployment (detailed analysis in Appendix A.1).
@@ 2.3 Related Work
Our work builds on three research threads: learned heuristics for constraint programming (gasse2019exact; cappart2021combinatorial), graph neural networks for structured optimization (velickovic2018graph; kipf2017semi), and reinforcement learning for combinatorial problems (sutton2018reinforcement; schulman2017proximal). Unlike prior work on search heuristics (gasse2019exact) or reactive scheduling (marques2012activity), MetaJuLS learns predictive propagation policies via meta-RL that generalize across languages and constraint types. A detailed survey appears in Appendix A.
@@ 3.1 Problem Formulation
The action space consists of all constraint propagators that are currently “dirty,” that is, constraints whose associated variables have experienced domain changes since their last propagation (bessiere2006constraint). At each time step , the agent selects an action , where is the subset of dirty constraints at time . The selected propagator is then executed, potentially reducing variable domains and marking additional constraints as dirty for future propagation (regin1994filtering; lecoutre2009str).
@@
where is the set of variables involved in constraint , is the current domain size of variable , and is a constraint-type-specific constant reflecting the complexity of the filtering algorithm. The constants reflect empirical operation counts: grammar rule constraints require checking two child domains (binary lookup), while lexical constraints require a single vocabulary check. These values are robust to perturbation; Appendix C.1 reports 2% performance variation for . For grammar rule constraints in parsing, we set (binary checks), while for lexical constraints (simpler lookups). For CP constraints, for binary constraints and for alldiff where is arity. This deterministic proxy avoids hardware-dependent wall-clock time measurements while accurately reflecting computational cost (schulman2017proximal). To validate the cost proxy, we measure correlation between proxy cost and wall-clock time across constraint types and sentence lengths. On PTB parsing, we find Pearson correlation (, ) between proxy cost and measured microseconds per propagator, confirming that optimizing the proxy improves real wall-clock time. The reward structure encourages the agent to discover policies that maximize search space reduction while minimizing solver overhead (debruyne1997optimal; gent2006watched).
@@ 4.1 Constituency Parsing: Penn Treebank
We compare against Berkeley Neural Parser (kitaev2018constituency), GPU-parallelized CKY (our reimplementation of klein2003accurate using batched CUDA kernels for span-parallel chart filling), A* search with neural heuristics (stern2017minimum), activity-based scheduling, VSIDS-style conflict-driven scheduling (prioritizing constraints that recently caused pruning), cost-normalized greedy (maximizing using our deterministic cost proxy), and supervised ranking (learning-to-rank constraints from state features, trained to imitate oracle best-first on development set). The meta-learned MetaJuLS policy is trained on PTB sections 2–21. Table 1 reports F1 scores and inference time. MetaJuLS achieves 1.6–1.9 speedups over GPU-parallelized CKY while maintaining within 0.2% F1 of Berkeley Neural Parser. MetaJuLS speedups increase with sentence length, achieving 1.91 on sentences tokens vs. 1.2 on short sentences (full scaling analysis in Appendix B.2).
@@ 4.2 Dependency Parsing: Universal Dependencies
We evaluate MetaJuLS on dependency parsing using Universal Dependencies (UD) v2.11 (zeman2022universal) across 10 diverse languages: English, Spanish, French, German, Chinese, Arabic, Finnish, Japanese, Russian, and Hindi. We implement a transition-based dependency parser where variables correspond to parser states, domains encode possible transitions, and constraints enforce arc consistency and well-formedness. All UD comparisons use the same dependency parser architecture, beam size, and batching; only the constraint scheduling policy changes, ensuring apples-to-apples comparison.
We compare against state-of-the-art dependency parsers (UDPipe (straka2016udpipe), Stanza (qi2020stanza)), activity-based scheduling, VSIDS-style conflict-driven scheduling, cost-normalized greedy, and MetaJuLS. The meta-learned MetaJuLS policy is trained on English UD data, then rapidly adapted to other languages with 5–10 gradient steps (requiring 5–15 seconds). Table 2 reports LAS (Labeled Attachment Score) and inference time. MetaJuLS achieves 1.5–1.8 speedups over activity-based scheduling while maintaining competitive LAS scores, as shown in Figure 2.
@@ 4.3 LLM Constrained Decoding
We evaluate on LogicBench (chen2024logicbench), which requires generating formal logical expressions that satisfy given premises. Variables correspond to token positions, domains contain logical operators and predicates, and constraints enforce well-formedness and logical consistency. To test whether MetaJuLS and speculative decoding are complementary, we perform a factorial experiment comparing: (1) baseline (neither method), (2) speculative decoding only, (3) MetaJuLS scheduling only, and (4) both methods combined. We compare against NeuroLogic Decoding (lu2021neurologic), Outlines (willard2023outlines), and LMQL (beurer2023lmql). MetaJuLS achieves 1.8 speedup over speculative decoding alone while maintaining 98.2% constraint satisfaction rate (vs. 94.1% for speculative decoding). When combined, speculative decoding + MetaJuLS achieves an additional 15% speedup over MetaJuLS alone, demonstrating orthogonality.
@@
Mechanistic analysis reveals the policy learns human-like easy-first parsing and discovers novel middle-out strategies for nested clauses (see Appendix B for detailed probing studies). Error analysis shows MetaJuLS reduces agreement errors by 38% and attachment errors by 22% compared to heuristic schedulers, though it underperforms on garden-path sentences (detailed breakdown in Appendix B.1).
@@ 4.5 Rapid Cross-Lingual and Cross-Task Adaptation
5 Generalization Beyond Language: CP Benchmarks
To test whether learned scheduling principles transfer beyond language, we evaluate on MiniZinc and XCSP constraint programming benchmarks. MetaJuLS achieves 6.6% average optimality gap vs. 8.5% for OR-Tools, with 0.63 normalized runtime and 94% solve rate (Table 5). Policies transfer bidirectionally: CP-trained policies achieve 89% of specialist performance on NLP tasks, and NLP-trained policies achieve 91% on CP benchmarks. On complex XCSP instances, MetaJuLS achieves 1.5 speedup over OR-Tools. Detailed per-category results and policy analysis appear in Appendix D.
| Solver | Avg. Gap (%) | Runtime (Norm.) | Solve Rate (%) |
| OR-Tools | 8.5 | 1.0 | 88% |
| JuLS (Base) | 9.1 | 1.1 | 84% |
| Activity-Based | 7.8 | 0.95 | 90% |
| MetaJuLS (Ours) | 6.6 | 0.63 | 94% |
6 Discussion and Conclusion
MetaJuLS shows that reinforcement learning can learn adaptive constraint-propagation schedules that outperform handcrafted and reactive heuristics by treating scheduling as a sequential decision problem with delayed, global effects (silver2016alphago; silver2017alphago; schulman2017proximal; sutton2018reinforcement; bessiere2006constraint; debruyne1997optimal; shihab2025detecting). Policy analysis indicates that the learned scheduler both rediscovers effective classical principles in simple regimes (e.g., Domain-over-Weight) and discovers non-trivial strategies in harder structured settings (e.g., prioritizing violation potential), supporting the claim that learning can capture and extend beyond standard heuristic design (apt2003principles; marriott1998programming; regin1994filtering; gent2006watched).
Empirically, MetaJuLS delivers 1.5–2.0 speedups over GPU-optimized baselines while staying within 0.2% of state-of-the-art accuracy on constituency parsing (PTB), dependency parsing (Universal Dependencies), and LLM constrained decoding (LogicBench, GSM8K-Constrained). Meta-learning enables rapid cross-lingual and cross-task adaptation in 5–10 gradient steps (5–15 seconds), avoiding hours of task-specific retraining. A safety-aware entropy-triggered fallback reduces worst-case accuracy gaps from 2.1% to 0.15% while preserving 1.4 average speedups. Finally, ablations against strong baselines (VSIDS-style, cost-normalized greedy, supervised ranking) show 1.2–1.5 speed improvements at matched accuracy, supporting the necessity of learned, state-aware scheduling for optimal structured inference.
7 Limitations
We evaluate on English parsing and constrained generation; the approach may require adaptation for morphologically rich languages or other structured prediction tasks. Meta-training requires 120 GPU-hours on 8 A100 GPUs (one-time cost), enabling rapid adaptation to new tasks with 5–10 gradient steps. The primary value proposition is latency reduction: MetaJuLS reduces p99 latency from 245ms to 142ms on constrained generation tasks, enabling deployment in latency-sensitive settings. We provide pre-trained checkpoints enabling deployment without retraining (detailed deployment ROI analysis in Appendix G). Our gains are relative to a specific CRF-based chart parser; modern neural parsers using different inference strategies may benefit differently. Casting a new NLP task as constraint propagation requires manual effort to define variables, domains, and constraints, though meta-learning reduces the adaptation burden once the constraint formulation is established.
While rapid adaptation achieves 85–92% of specialist performance in many cases, some domains transfer less effectively (e.g., TSP to Scheduling: 74% of specialist performance), indicating domain-specific knowledge remains important. The current approach assumes deterministic propagators; extending to stochastic or approximate propagators is future work (detailed transfer analysis and robustness discussion in Appendix G.2).
Acknowledgment and Reproducibiity
We utilized AI assistance for grammar corrections and a few plot-related purposes. Everything was validated by the authors.
Code, trained models, and experiment scripts will be available upon acceptance. We include: (1) Full training and evaluation code; (2) Pre-trained meta-learned policy checkpoints; (3) Scripts to reproduce all tables and figures; (4) Preprocessed data splits (PTB, UD, and LLM datasets). Meta-training requires approximately 120 GPU-hours on 8 A100 GPUs (one-time cost), enabling rapid adaptation to new tasks with 5–10 gradient steps (5–15 seconds on a single GPU). Inference adds 3% overhead to base parser/LLM runtime. All experiments use identical hardware (A100 GPUs), batch sizes (1 for parsing, 8 for constrained decoding), and measurement protocols (warmup runs excluded, median of 5 runs).
Appendix A Extended Related Work
Activity-based heuristics, inspired by the VSIDS (Variable State Independent Decaying Sum) strategy from SAT solving (moskewicz2001chaff; marques2012activity), represent a significant advance over static scheduling (regin1994filtering; lecoutre2009str). These methods maintain activity scores for constraints based on their historical contribution to search progress, dynamically prioritizing constraints that have recently caused domain reductions or failures (audemard2013glucose; debruyne1997optimal). However, activity-based heuristics remain reactive rather than predictive, adjusting priorities only after observing constraint behavior rather than anticipating which constraints will be most effective in the current context.
Our work builds on these foundations by combining graph neural networks with reinforcement learning to learn adaptive propagation policies (velickovic2018graph; schulman2017proximal). Unlike previous approaches that rely on handcrafted features or reactive heuristics (wallace1996practical; marques2012activity), MetaJuLS learns to extract relevant information from the constraint-variable graph structure itself (kipf2017semi; hamilton2017inductive), enabling the discovery of novel scheduling strategies that adapt to both problem structure and solver state (bessiere2006constraint).
A.1 Why RL Is Necessary: Detailed Analysis
Appendix B Mechanistic Analysis and Policy Interpretability